### v1

@@ 4.2 Posterior Predictions
Figure 6 presents a detailed view of the posterior predictive distributions, $\pi(\widehat{Y}^{C}_{s,t})$ and $\pi(\widehat{Y}^{B}_{s,t})$, obtained from the second stage model. To assess predictive performance, we conduct posterior predictive checks based on the threshold exceedance probabilities of fire count and burnt area within the test set. The empirical exceedance probabilities at a given threshold $h$ are defined as
$\widehat{P}(Y^{C}>h)=\frac{1}{|\mathcal{I}|}\sum_{(s,t)\in\mathcal{I}}\mathbbm{1}(Y^{C}_{s,t}>h),\quad\widehat{P}(Y^{B}>h)=\frac{1}{|\mathcal{I}|}\sum_{(s,t)\in\mathcal{I}}\mathbbm{1}(Y^{B}_{s,t}>h),$
where $\mathcal{I}$ denotes the set of all spatial and temporal indices in the test set, and $|\mathcal{I}|$ is its cardinality. These quantities represent the overall proportions of exceedances in the test set, aggregated over space and time, and are not conditioned on specific locations or time points. The upper panels of Figure 6 show the posterior predictive check of the exceedance probabilities over various thresholds based on 1000 posterior predictive replicates. Uncertainty is notably higher at lower thresholds (e.g., 5 fires or 10 hectares) and diminishes as the threshold increases. Although the empirical exceedance probabilities (red dashed lines) may deviate from the centres of the predictive distributions, they consistently fall within the 50th to 90th percentiles, suggesting satisfactory model calibration.
@@ 4.3.1 XGBoost model interpretation
The XGBoost model provides point forecasts of wildfire activity, while the latent Gaussian model primarily quantifies associated uncertainty. It is therefore essential to understand which covariates most strongly influence the predictions generated by XGBoost. To this end, we assess covariate importance using SHapley Additive exPlanations (SHAP) values (Lundberg and Lee,, 2017).
SHAP values offer a principled approach to attributing the marginal contribution of each covariate to the model output, drawing on the concept of Shapley values from cooperative game theory (Shapley,, 1953). For a model $f$ fitted on a covariate set $M=\{\widetilde{x}_{1},\widetilde{x}_{2},\cdots,\widetilde{x}_{m}\}$, and a subset $S\subseteq M$, the SHAP value $\phi_{j}$ of covariate $\widetilde{x}_{j}$ is the average difference between the predictions $f(S\cup\{\widetilde{x}_{j}\})$ and $f(S)$ over all possible $S$. Formally, $\phi_{j}$ is defined by
@@
Since $f$ typically requires the full covariate set $M$, the output for a reduced set $S$ is approximated as $f(S)=\mathbb{E}[f(M)\mid S]$. For tree-based models such as XGBoost, this conditional expectation can be efficiently computed using the algorithm proposed by Lundberg et al., (2018).
@@
Lundberg and Lee, (2017) showed that (13) is the unique additive representation that satisfies local accuracy, missingness, and consistency. This decomposition provides an intuitive and theoretically grounded measure of covariate influence, based on both the sign and magnitude of the SHAP values.
Figure 7 displays the SHAP values for the ten most influential covariates in the XGBoost models for fire count (5.3) and burnt area (2). In both models, autoregressive terms dominate the set of top covariates, suggesting that historical wildfire activity contributes more to predictive accuracy than the environmental variables. The most influential covariates are the averages of fire count and burnt area of the three months centred at the forecast month over the past three years, aggregated at the council level (conc_fc_hist_3 and conc_ba_hist_3, respectively). While high values of these autoregressive covariates do not always lead to large forecasts, they are generally positively correlated with wildfire events. Among the environmental covariates, average air temperature (Temp) and relative humidity (RHumi) at the monthly council level have the largest mean absolute SHAP values. Notably, the SHAP values of these variables exhibit an intuitive non-causal relationship with the wildfire activity. For instance, high fire count and burnt area are often associated with elevated temperatures (reflected by large positive SHAP values), whereas low fire activity can occur across a broader range of temperature levels (indicated by the mix of blue and red at the lower end of the SHAP value), consistent with domain expectations.
@@ 4.3.2 Latent Gaussian effect of Year and Covariates
Figure 8 shows the posterior estimates of the year-specific effects $T(\cdot)$ and the effects of XGBoost predictions $R(\cdot)$ in $\eta^{C}$ and $\eta^{B}$ as specified in (7). While the year effects show some variation across time, only the effect for 2017 in the fire presence model is significantly greater than zero. This is consistent with the historical record, as Portugal experienced the highest number of fire ignitions and the largest total burnt area in 2017 over the past decade. Given that the XGBoost model already incorporates temporal information, the lack of significant year effects in most other years suggests that the XGBoost predictions capture the interannual variation effectively.
@@ 4.3.3 Shared Spatio-Temporal Effects
Figure 9: Posterior mean of the council-level spatio-temporal effect $G_{c}(\cdot)$ by grouped month (left), and average of the posterior mean of the district-level spatio-temporal effect $G_{d}(\cdot)$ (right).
Figure 9 displays the posterior means of the council-level spatio-temporal effect $G_{c}(\cdot)$, grouped by month, and the average district-level spatio-temporal effect $G_{d}(\cdot)$ across all time indices. During the high-risk wildfire season, particularly in July, August, and October, the council-level effects exhibit greater spatial variability, with notable contrasts between neighbouring councils. In contrast, during the remaining months, the spatial effects are more homogeneous and show smooth transitions across adjacent regions. Noteworthy exceptions include Montalegre and Vinhais councils: the former shows unusually elevated effects in January and March, while the latter displays pronounced effects in April relative to its surrounding areas. At the district level, a general spatial gradient is evident, with higher effect values observed in the southwestern districts and lower values in the northeast. However, the scale of the average district level is pretty small, with a maximum value around 5% of the maximum of the council level effect.
The estimated posterior means of the scaling parameters for the shared effects are $\widehat{\beta}_{1}^{C}=0.78$, $\widehat{\beta}_{1}^{B}=0.41$, $\widehat{\beta}_{2}^{C}=0.56$, $\widehat{\beta}_{2}^{B}=0.31$. The scaling parameters associated with fire count are consistently larger than those for burnt area, indicating that the shared spatio-temporal effects for fire count are more closely aligned with those for fire presence. This finding is intuitive, as fire count is derived from aggregating fire presence events, whereas burnt area is a continuous measure associated with those events.
@@ 4.4 Posterior Distributions of eGP Parameters
We now provide insights into the posterior inference for the parameters $\xi$ and $\kappa$ from the perspective of implementing the eGP likelihood within the INLA framework. These parameters are treated as global hyperparameters in the latent Gaussian model, meaning they are shared across all observations. Figure 10 shows the prior and posterior distributions for $\xi$ and $\kappa$. The posterior of $\xi$, compared to its prior centred around zero, shifts markedly to the right and concentrates near $0.4$. This suggests that the fitted eGP likelihood possesses a heavy-tailed structure, which may be suitable for capturing the extreme behaviour observed in burnt area data. Notably, the mode of the posterior for $\xi$ lies close to the upper bound of its prior, potentially conflicting with the imposed constraint of (0.5,0.5)(-0.5,0.5). However, we argue that the posterior of $\xi$ is particularly sensitive to the skewness and overall shape of the response distribution, and that its value has only a minor influence on the posterior predictive distribution of burnt area. This argument is further examined in Section 5.2. Therefore, rather than focusing on the interpretability of the posterior of $\xi$, it is more critical to ensure that the fitted eGP likelihood retains desirable properties, such as finite variance, by appropriately constraining the prior. As for $\kappa$, its posterior distribution deviates substantially from the prior, with mode concentrated around $4.7$, though the posterior of $\kappa$ shows larger dispersion compared to the posterior of $\xi$. This indicates that the posterior predictive probability $f_{\text{eGP}}(y_{s,t}\mid\eta_{s,t}^{B},\xi,\kappa)$ does not have a singularity at 0, and its shape is approximately bell-like, though potentially skewed (see Figure 11).
@@ 5.1 Data Transformation and $\xi$ in eGP
The square root transformation strikes a practical balance, moderating skewness without unduly suppressing large values. Our empirical investigations, which are guided by posterior predictive checks and sensitivity analyses, show that this transformation yields stable and interpretable posteriors for $\xi$, consistent with prior beliefs and with tail behaviour observed in historical burnt area data. This choice aligns with modelling choices in recent wildfire literature, such as Cisneros et al., (2024). While alternative power transformations (e.g., cube root, fourth root, or logarithm) influence the posterior of $\xi$, they result in nearly identical posterior predictive distributions once back-transformed, reinforcing the square root as a pragmatic and robust choice.
@@ 5.2 Similar Performance to Gamma and Weibull Likelihoods
Somewhat unexpectedly, the eGP likelihood does not significantly outperform the Gamma or Weibull likelihoods, despite the fact that the burnt area data are clearly heavy-tailed and the eGP distribution is explicitly designed to accommodate such behaviour. While the Gamma distribution is light-tailed, the Weibull distribution, given the posterior mean of its shape parameter at approximately 1.39, effectively exhibits heavy-tailed behaviour in this context. Figure 11 compares the three likelihoods using estimated hyperparameters. The linear predictors have been adjusted so that the medians of the resulting distributions align at approximately 0.8, facilitating a shape comparison. Although the eGP distribution shows a heavier tail, the overall shapes of the three densities are relatively similar. These findings are consistent with the results presented in Table (1), which indicate that the data are largely insensitive to the tail properties of the assumed likelihood.
These results suggest that extreme value theory does not offer a universally superior solution for modelling extremes. One plausible reason lies in the hierarchical structure of the model. Within this framework, observations are assumed conditionally independent given their respective linear predictors, which are functions of Gaussian latent effects. The central tendency and quantiles of the marginal distribution (e.g., its mean or $\alpha$-quantile) are controlled by the linear predictor. Since hyperparameters such as $\xi$ and $\kappa$ are shared across all observations, their influence on individual predictive densities is limited relative to the localised effect of the linear predictor. When the latent structure is sufficiently expressive, especially through the inclusion of informative covariates such as the XGBoost predictions, the linear predictor can effectively account for both moderate and extreme observations. Consequently, the observations tend to cluster within the high-density region of the fitted marginal distribution, diminishing the role of the tail parameter in shaping the likelihood. As a result, the contribution of tail behaviour to overall model performance becomes less critical, which explains the comparable predictive accuracy across the three likelihoods.
Figure 11: Densities of Gamma (shape = 1.87), Weibull (shape = 1.39) and eGP ($\kappa=4.72,\,\xi=0.42$) likelihoods, based on estimated hyperparameters. The densities are scaled via the linear predictor to have approximately equal medians (around 0.8).
@@ 5.3 Longer Forecasting Horizons
The current forecasting horizon considered in our framework is one month—that is, we generate forecasts $y_{t+1}$ conditional on the past observations $y_{1:t}$ and additional covariates $\widetilde{\boldsymbol{x}}_{t}$. Extending this to longer-term forecasts (e.g. forecasting $y_{t+h}$ for $h>1$) presents a key challenge: neither the XGBoost model nor the latent Gaussian model is inherently designed to produce multi-horizon predictions in a unified structure. As such, it is not straightforward to generate forecasts across multiple future time points using a single model fit.
A practical solution is to apply the full two-stage modelling framework (as illustrated in Figure 3) independently for each forecasting horizon $h$. While the process has been detailed for $h=1$, the extension to $h=2,3,\ldots$ involves adjusting and retraining the XGBoost model to predict future outcomes based on features lagged by $h$ months. Specifically, the XGBoost models can be modified to produce:

### v2

@@ 4.2 Posterior Predictions
Figure 7 presents a detailed view of the posterior predictive distributions, $\pi(\widehat{Y}^{C}_{s,t+1})$ and $\pi(\widehat{Y}^{B}_{s,t+1})$, obtained from the second stage model. To assess predictive performance, we conduct posterior predictive checks based on the threshold exceedance probabilities of fire count and burnt area within the test set. The empirical exceedance probabilities at a given threshold $h$ are defined as
$\widehat{P}(Y^{C}>h)=\frac{1}{|\mathcal{I}|}\sum_{(s,t+1)\in\mathcal{I}}\mathbbm{1}(Y^{C}_{s,t+1}>h),\quad\widehat{P}(Y^{B}>h)=\frac{1}{|\mathcal{I}|}\sum_{(s,t+1)\in\mathcal{I}}\mathbbm{1}(Y^{B}_{s,t+1}>h),$
where $\mathcal{I}$ denotes the set of all spatial and temporal indices in the test set, and $|\mathcal{I}|$ is its cardinality. These quantities represent the overall proportions of exceedances in the test set, aggregated over space and time, and are not conditioned on specific locations or time points.
The upper panels of Figure 7 show the posterior predictive check of the exceedance probabilities over various thresholds based on 1000 posterior predictive replicates. Uncertainty is notably higher at lower thresholds (e.g., 5 fires or 10 hectares) and diminishes as the threshold increases. When compared with empirical exceedance probabilities, the model performs well for burnt area: the observed values (red dashed lines) consistently fall within the 50th–90th percentiles of the posterior predictive distribution. For fire counts, some deviations occur at lower thresholds, but the observed statistics do not systematically fall outside the predictive envelopes.
@@ 4.3.1 XGBoost model interpretation
The XGBoost model provides point forecasts of wildfire activity, while the latent Gaussian model primarily quantifies associated uncertainty. It is therefore essential to understand which covariates most strongly influence the predictions generated by XGBoost. To this end, we assess covariate importance using SHapley Additive exPlanations (SHAP) values (Lundberg and Lee, 2017).
SHAP values offer a principled approach to attributing the marginal contribution of each covariate to the model output, drawing on the concept of Shapley values from cooperative game theory (Shapley, 1953). For a model $f$ fitted on a covariate set $M=\{\widetilde{x}_{1},\widetilde{x}_{2},\cdots,\widetilde{x}_{m}\}$, and a subset $S\subseteq M$, the SHAP value $\phi_{j}$ of covariate $\widetilde{x}_{j}$ is the average difference between the predictions $f(S\cup\{\widetilde{x}_{j}\})$ and $f(S)$ over all possible $S$. Formally, $\phi_{j}$ is defined by
@@
Since $f$ typically requires the full covariate set $M$, the output for a reduced set $S$ is approximated as $f(S)=\mathbb{E}[f(M)\mid S]$. For tree-based models such as XGBoost, this conditional expectation can be efficiently computed using the algorithm proposed by Lundberg et al. (2018).
@@
Lundberg and Lee (2017) showed that (13) is the unique additive representation that satisfies local accuracy, missingness, and consistency. This decomposition provides an intuitive and theoretically grounded measure of covariate influence, based on both the sign and magnitude of the SHAP values.
Figure 8 displays the SHAP values for the ten most influential covariates in the XGBoost models for fire count (5.3) and burnt area (2). In both models, autoregressive terms dominate the set of top covariates, suggesting that historical wildfire activity contributes more to predictive accuracy than the environmental variables. The most influential covariates are the averages of fire count and burnt area of the three months centred at the forecast month over the past three years, aggregated at the council level (conc_fc_hist_3 and conc_ba_hist_3, respectively). While high values of these autoregressive covariates do not always lead to large forecasts, they are generally positively correlated with wildfire events. Among the environmental covariates, average air temperature (Temp) contributes most strongly to the fire count model, while both Temp and relative humidity (RHumi) are most important for the burnt area model. The SHAP values of these variables exhibit an intuitive, non-causal relationship with the wildfire activity. For instance, high fire count and burnt area are often associated with elevated temperatures (reflected by large positive SHAP values), whereas low fire activity can occur across a broader range of temperature levels (indicated by the mix of blue and red at the lower end of the SHAP value), consistent with domain expectations.
@@ 4.3.2 Latent Gaussian effect of Year and Covariates
Figure 9 shows the posterior estimates of the year-specific effects $T(\cdot)$ and the effects of XGBoost predictions $R(\cdot)$ in $\eta^{C}$ and $\eta^{B}$ as specified in (7). While the year effects vary over time, only those for fire presence in 2017 and 2018 are significantly different from zero. This aligns with historical records: Portugal experienced the highest number of fire ignitions in 2017 in the past decade, followed by a sharp decline in 2018, possibly due to fire regulation adjustments after the catastrophic wildfire activity of 2017. As the XGBoost model already incorporates temporal information, the absence of significant year effects in most other years suggests that its predictions capture interannual variation effectively.
@@ 4.3.3 Shared Spatio-Temporal Effects
Figure 10: Posterior mean of the council-level spatio-temporal effect $G_{c}(\cdot)$ by grouped month (left), and average of the posterior mean of the district-level spatio-temporal effect $G_{d}(\cdot)$ (right), aggregated over all time indices.
Figure 10 displays the posterior means of the council-level spatio-temporal effect $G_{c}(\cdot)$, grouped by month, and the average district-level spatio-temporal effect $G_{d}(\cdot)$ across all time indices. Their uncertainty, quantified by 0.025 and 0.975 posterior quantiles, is provided in Figure A.2. During the high-risk wildfire season, particularly in July, August, and October, the council-level effects exhibit greater spatial variability, with notable contrasts between neighbouring councils. In contrast, during the remaining months, the spatial effects are more homogeneous and show smooth transitions across adjacent regions. Noteworthy exceptions include Montalegre and Vinhais councils: the former shows unusually elevated effects in January and March, while the latter displays pronounced effects in April relative to its surrounding areas. At the district level, a general spatial gradient is evident, with higher effect values observed in the northern districts and lower values in the south. However, the scale of the average district level is relatively small, with the maximum magnitude reaching only about 15% of that in the council-level.
The estimated posterior means of the scaling parameters for the shared effects are $\widehat{\beta}_{1}^{C}=-0.69$, $\widehat{\beta}_{1}^{B}=-0.43$, $\widehat{\beta}_{2}^{C}=0.77$, $\widehat{\beta}_{2}^{B}=0.27$. The signs of these parameters are not straightforward to interpret, given the dependence between council- and district-level spatio-temporal effects. Focusing instead on magnitudes within each wildfire quantity, the scaling parameters for fire count are consistently larger than those for burnt area, suggesting that the shared spatio-temporal effects for fire count are more strongly correlated with those for fire presence. This is because fire count is obtained by aggregating fire presence events, whereas burnt area is a continuous measure conditional on those events.
@@ 4.4 Posterior Distributions of eGP Parameters
We now provide insights into the posterior inference for the parameters $\xi$ and $\kappa$ from the perspective of implementing the eGP likelihood within the INLA framework. These parameters are treated as global hyperparameters in the latent Gaussian model, meaning they are shared across all observations. Figure 11 shows the prior and posterior distributions for $\xi$ and $\kappa$. The posterior of $\xi$, compared to its prior centred around zero, shifts markedly to the right and concentrates near $0.45$. This suggests that the fitted eGP likelihood possesses a heavy-tailed structure, which may be suitable for capturing the extreme behaviour observed in burnt area data. Notably, the mode of the posterior for $\xi$ lies close to the upper bound of its prior, potentially conflicting with the imposed constraint of (0.5,0.5)(-0.5,0.5). However, we argue that the posterior of $\xi$ is particularly sensitive to the skewness and overall shape of the response distribution, and that its value has only a minor influence on the posterior predictive distribution of burnt area. This argument is further examined in Section 5.2. Therefore, rather than focusing on the interpretability of the posterior of $\xi$, it is more critical to ensure that the fitted eGP likelihood retains desirable properties, such as finite variance, by appropriately constraining the prior. As for $\kappa$, its posterior distribution deviates substantially from the prior, with mode concentrated around $4.6$, though the posterior of $\kappa$ shows larger dispersion compared to the posterior of $\xi$. This indicates that the posterior predictive probability $f_{\text{eGP}}(y_{s,t+1}\mid\eta_{s,t+1}^{B},\xi,\kappa)$ does not have a singularity at 0, and its shape is approximately bell-like, though potentially skewed (see Figure 12).
@@ 5.1 Data Transformation and $\xi$ in eGP
The square root transformation strikes a practical balance, moderating skewness without unduly suppressing large values. Our empirical investigations, which are guided by posterior predictive checks and sensitivity analyses, show that this transformation yields stable and interpretable posteriors for $\xi$, consistent with prior beliefs and with tail behaviour observed in historical burnt area data. This choice aligns with modelling choices in recent wildfire literature, such as Cisneros et al. (2024). While alternative power transformations (e.g., cube root, fourth root, or logarithm) influence the posterior of $\xi$, they result in nearly identical posterior predictive distributions once back-transformed, reinforcing the square root as a pragmatic and robust choice.
@@ 5.2 Similar Performance to Gamma and Weibull Likelihoods
Although the eGP likelihood outperforms the Gamma and Weibull likelihoods in Table 1, the differences are modest, particularly relative to the Weibull, which also exhibits heavy-tailed behaviour with a posterior mean shape parameter of approximately 1.35. Figure 12 compares the three likelihoods using estimated hyperparameters. The linear predictors have been adjusted so that the medians of the resulting distributions align at approximately 0.8, facilitating a shape comparison. While the eGP distribution displays a heavier tail than the other two, the overall shapes of the three densities are broadly similar, with the largest differences appearing near the mode. These findings, consistent with Table 1, indicate that refining the likelihood to capture tail behaviour has limited impact on forecast performance; instead, accuracy is primarily determined by the flexibility of the latent structure and the incorporation of informative covariates into the linear predictor. Indeed, within the hierarchical structure of our model, observations are assumed conditionally independent given their respective linear predictors, which are functions of Gaussian latent effects. The central tendency and quantiles of the marginal distribution (e.g., its mean or $\alpha$-quantile) are controlled by the linear predictor. Since hyperparameters such as $\xi$ and $\kappa$ are shared across all observations, their influence on individual predictive densities is limited relative to the localised effect of the linear predictor. When the latent structure is sufficiently expressive, especially through the inclusion of informative covariates such as the XGBoost predictions, the linear predictor can effectively account for both moderate and extreme observations. Consequently, the observations tend to cluster within the high-density region of the fitted marginal distribution, diminishing the role of the tail parameter in shaping the likelihood. As a result, the contribution of tail behaviour to overall model performance becomes less critical, which explains the comparable predictive accuracy across the three likelihoods.
Figure 12: Densities of Gamma (shape = 1.80), Weibull (shape = 1.35) and eGP ($\kappa=4.64,\,\xi=0.45$) likelihoods, based on estimated hyperparameters. The densities are scaled via the linear predictor to have approximately equal medians (around 0.8).
@@ 5.3 Longer Forecasting Horizons
The current forecasting horizon considered in our framework is one month. That is, we generate forecasts $y_{t+1}$ conditional on the past observations $y_{1:t}$ and additional covariates $\widetilde{\boldsymbol{x}}_{t}$. Extending this to longer-term forecasts (e.g. forecasting $y_{t+h}$ for $h>1$) presents a key challenge: neither the XGBoost model nor the latent Gaussian model is inherently designed to produce multi-horizon predictions in a unified structure. As such, it is not straightforward to generate forecasts across multiple future time points using a single model fit.
A practical solution is to apply the full two-stage modelling framework (as illustrated in Figure 3) independently for each forecasting horizon $h$. While the process has been detailed for $h=1$, the extension to $h=2,3,\ldots$ involves adjusting and retraining the XGBoost model to predict future outcomes based on features lagged by $h$ months. We emphasise that this approach does not constitute a fully coherent joint multi-horizon forecasting model, but rather a pragmatic, horizon-specific solution that remains compatible with the INLA framework and preserves interpretability and uncertainty quantification. Specifically, the XGBoost models can be modified to produce:
@@
Since the most influential covariates in the XGBoost models for $h=1$ are derived from historical wildfire activity at the same forecast month (See Figure 8 and Table A.2), these predictors tend to continue to dominate the feature set when $h>1$. Consequently, the predictive accuracy of the XGBoost models does not deteriorate substantially for longer forecast horizons ($h>1$). This stability implies that the proposed multi-horizon forecast method is less susceptible to error propagation over time and can deliver robust uncertainty quantification. Table 2 presents a comprehensive comparison of forecast accuracy and uncertainty for model M1 with the eGP likelihood across horizons $h=1,2,3$. In Stage 1, mean squared error (MSE) is used to assess the performance of the XGBoost models. In Stage 2, forecast accuracy is evaluated using the same metrics as in Table 1, while uncertainty is quantified by the average length of the 95% credible interval derived from INLA. Only minor variations in predictive performance and uncertainty are observed across horizons from one to three months, with no evidence of a systematic degradation in performance.
Table 2: Comparison of prediction accuracy and uncertainty for model M1 across forecast horizons $h=1,2,3$. Mean squared error (MSE) evaluates the accuracy of the two Stage 1 XGBoost models. Stage 2 prediction metrics are identical to those reported in Table 1. Uncertainty in Stage 2 is quantified by the average length of the 95% credible interval.
Forecast Horizon h=1 h=2 h=3
Stage 1 Prediction MSE for count 0.34 0.36 0.30
MSE for burnt area 8.26 8.27 8.24
Stage 2 Prediction AUC 0.862 0.857 0.864
CRPS 4.75 4.61 4.61
Unweighted $r^{C}$ 362 367 345
Weighted $r^{C}$ 4.84 5.11 4.83
Unweighted $r^{B}$ 565 585 599
Weighted $r^{B}$ 14.2 14.8 14.4
Stage 2 Uncertatinty Avg. CI for count 3.38 3.49 3.50
Avg. CI for burnt area 24.33 24.28 23.92
5.4 Choices of the Model in Stage 1
In this study, XGBoost is adopted in Stage 1 to generate one-month-ahead wildfire forecasts at the council level. Since Stage 1 only requires point forecasts, XGBoost is not the only viable choice. Alternative models, such as ensemble methods (e.g., Random Forest; Breiman, 2001) or deep learning approaches for time series forecasting (e.g., long short-term memory networks; Hochreiter and Schmidhuber, 1997, and attention-based models; Lim et al., 2021; Zhou et al., 2021), could serve a similar role. In principle, multiple models targeting the same response could be fitted in Stage 1, with their outputs jointly incorporated as covariates in Stage 2. This strategy is not pursued here because the number of hyperparameters associated with existing model components is already close to 20, which is near the recommended upper limit for INLA. When selecting a single model, numerous studies have demonstrated that XGBoost achieves state-of-the-art performance in both time series forecasting (Elsayed et al., 2021; Januschowski et al., 2022; Santoro et al., 2024) and tabular data modelling (Gorishniy et al., 2021; Grinsztajn et al., 2022). For these reasons, XGBoost is chosen as the Stage 1 model in our framework.